OpenStudy (anonymous):

What is the common ratio of the geometric sequence 16, 24, 36, 54, ...? 2/3 3/2 1 8

OpenStudy (anonymous):

@twvogels

OpenStudy (solomonzelman):

divide any term by the term before it to find the ratio.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

why are you tagging me?

OpenStudy (anonymous):

We can find out by dividing each term by the one before it:\[\frac{54}{36}=1.5\]We can check another one to be sure\[\frac{36}{24}=1.5\]

OpenStudy (anonymous):

Or you can leave it as a fraction\[\frac{3}{2} = 1\frac{1}{2}\]

OpenStudy (anonymous):

ok thx